OpenStudy (anonymous):

PLease help Point R is located at (−5, 1), and S is located at (4, −6). Find the x value of the point that lies halfway between R and S. -2 -3 -0.5 -4.5

OpenStudy (anonymous):

\[\frac{x1 + x2}{2} , \frac {y1 + y2}{2}\]

OpenStudy (anonymous):

Can you help me apply it. I dont understand this

OpenStudy (anonymous):

sure so your point is (-5,1) that is (x1,y1) and the point (4,-6) is (x2,y2)

OpenStudy (anonymous):

\[\frac{-5 + 4}{2}, \frac{-6 + 1}{2}\]

OpenStudy (anonymous):

- 0.5, -2.5 thats what I got..

OpenStudy (anonymous):

I got -3. am I correct.

OpenStudy (anonymous):

you dont add them. it just wants the x value which is the first one

OpenStudy (anonymous):

so -0.5 Im really confused.

OpenStudy (anonymous):

@Austin6i6